OpenStudy (anonymous):

find slope intercept equation for the line with the indicated slope and y-intercept slope7, y-intercept (0,0)

OpenStudy (anonymous):

y-int = 0=b slope= 7=m y=mx+b

OpenStudy (anonymous):

plug the values given and you got yourself an equation buddy

OpenStudy (waleed_imtiaz):

It mean that y-intercept. c=0..... and m(slope)=7..... so put the values in the equation... y=mx+c.... y=7x or 7x-y=0 is the equation

OpenStudy (anonymous):

thanks for your help